OpenStudy (anonymous):

A ceiling fan has 17 -inch blades (so the radius of the circular fan is 17 inches). Suppose the linear speed of the tip of a blade is 9 feet per second. (a) Find the angular speed of the fan in radians per minute. (b) Find the number of revolutions a blade makes per minute. Do not round any intermediate computations, and round your answer to the nearest whole number.

OpenStudy (anonymous):

@Brrandyn @Compassionate

OpenStudy (anonymous):

i don't know if its 17 or 3 for a and 61 or 249 for b or if I'm totally wrong...

OpenStudy (brrandyn):

I've drawn the given scenario (basic model): Provided that the linear speed (of the blade "tip") is 9, use the two following equations to find the solution. Angular Velocity: \[\omega =\frac{ \theta }{ t }\] Linear Velocity: \[\ V=\frac{ \ r*θ }{t}\] b. Replace "V" with provided value and solve. \[V=\frac{ 17*(9*2\pi)}{1s}=480 ft/sec\] Gosh, I'm sorry for the slow response time, drawing panel refuses to close and unable to type anything when using the equation window! Very infuriating, haha.

OpenStudy (brrandyn):

Try solving for the angular velocity (: .